how many gallons of 70% antifreeze should be mixed with 10 gallons of 30% antifreeze to obtain 45% antifreeze.
Answer by mananth(16946)

------ percent ---------------- quantity
Antifreez1 I 80---------------- 40 gallons
Antifreeze II 10----------------x gallons
Total 30 ----------------40+x gallons
30*10+70*x=45(10+x)
300+70x=450+45x
70x-45x=450-300
25x=150
/25
x=6 gallons 70 % Antifreeze II